What would be the result if you combine 33 grams of methane with 289 grams of oxygen


1
Expert's answer
2021-02-16T07:46:53-0500

if you combine 33 grams of methane with 289 grams of oxygen you get water and carbon dioxide


The equation for the reaction is:

CH4 +2O2 → CO2 + 2H2O

IF 16 g CH4 +64 g O2 → 44 g CO2 + 36 g H2O

33g CH4 + 289g O2 will produce ??

16 g Methane produce 36 g water

Therefore, 33g CH4 will produce?

33 x 36/16= 74.25g H2O

CH4=16g

12g is carbon

therefore 33g of CH4 contains 33cx 12/16= 24.75 carbon

each mole of carbon (12g) burns to 44g of C02

Therefore, 24.75g carbon x 44/12 =90.75g of CO2

33g CH4 + 289g O2 → 90.75g CO2 + 74.25g H2O